# Distance of Quadratic Algebraic Numbers from the Middle-Third Cantor Set

### v1

@@ Abstract
We study the distance from quadratic irrational numbers to the middle-third Cantor set $C$. Mahler asked whether $C$ contains any irrational algebraic numbers; this remains open even for quadratic irrationals. Rather than assuming an answer to this problem, we obtain uniform lower bounds for the distance from a quadratic irrational $\alpha$ to $C$ in terms of the height $H$ of the minimal polynomial of $\alpha$.
We encode $\alpha$ by its orbit under the map $x\mapsto 3x\bmod 1$ and define the exit time $\operatorname{exit}(\alpha)$ as the first iterate that enters the middle interval $[1/3,2/3]$. Our main unconditional result is a quadratic exit bound $\operatorname{exit}(\alpha)\leq A(\log_{3}H)^{2}+B$ for absolute constants $A,B>0$, valid for all quadratic irrationals whose orbit stays a fixed small distance away from the coarse Cantor boundaries. As a consequence we obtain a distance lower bound $\operatorname{dist}(\alpha,C)\geq H^{-\kappa\log H}$ for some constant $\kappa>0$.
On the dynamical side we classify orbits by an $L/M/R$ coding and prove that the total number of visits to the right interval $[2/3,1)$ is $O(\log H)$. A finite case analysis on a bounded portion of the orbit is reduced to checking a finite list of explicit affine inequalities on subintervals of $[0,1]$, which we verify with short computer scripts; all Diophantine and dynamical estimates are proved by hand.
MSC 2020: 11J04 (primary), 28A80
Keywords: Cantor set, algebraic numbers, Mahler’s problem, Diophantine approximation, quadratic irrationals
@@ 1 Introduction
Let $\mathcal{C}\subset[0,1]$ denote the standard middle-third Cantor set. It is a classical object of fractal geometry, of Hausdorff dimension $\log 2/\log 3$, arising by repeatedly removing the open middle third from each closed interval. In base $3$ it can be described as
$\mathcal{C}=\Big\{\theta\in[0,1]:\theta=0.\varepsilon_{1}\varepsilon_{2}\cdots\text{ in base }3\text{ with }\varepsilon_{n}\in\{0,2\}\ \forall n\Big\}.$
It is a long-standing open problem, going back to Mahler, to understand the algebraic points of $\mathcal{C}$. In his 1984 paper on Diophantine approximation on Cantor-like sets Mahler asked whether $\mathcal{C}$ contains any irrational algebraic numbers at all [3]. This remains wide open: in particular, it is not known whether a quadratic irrational can lie in $\mathcal{C}$. Levesley–Salp–Velani explicitly note that, to the best of current knowledge, it is not even known whether or not quadratic irrationals avoid the middle-third Cantor set [4, Remark 1.7]. Our results therefore do not presuppose that algebraic points are excluded from $\mathcal{C}$; rather, we obtain quantitative lower bounds for the distance from $\mathcal{C}$ that hold uniformly for all quadratic irrationals whose orbits stay a fixed distance away from the coarse Cantor boundaries.
1.1 Exit time and dynamical formulation
Let $\alpha\in\mathbb{R}$ and consider the orbit
$\theta_{N}\coloneqq\{3^{N}\alpha\}\in[0,1],\qquad N=0,1,2,\dots.$
We partition $[0,1]$ into three intervals
$L\coloneqq[0,1/3),\quad M\coloneqq[1/3,2/3),\quad R\coloneqq[2/3,1).$
The Cantor set $\mathcal{C}$ is the set of all $\theta_{0}$ whose orbit under the map $\theta\mapsto\{3\theta\}$ never enters $M$. Thus for $\theta_{0}\notin\mathcal{C}$ the first time we enter $M$ is a natural dynamical measure of how far $\theta_{0}$ is from $\mathcal{C}$.
@@ 1.1 Exit time and dynamical formulation
Let $\alpha\in\mathbb{R}$. We define the exit time
$\mathrm{exit}(\alpha)\coloneqq\inf\{N\geq 1:\theta_{N-1}\in M\}\in\mathbb{N}\cup\{\infty\}$
with the convention that $\mathrm{exit}(\alpha)=\infty$ if the orbit never enters $M$. When $\mathrm{exit}(\alpha)<\infty$ we say that $\alpha$ exits by time $N=\mathrm{exit}(\alpha)$.
Thus $\mathrm{exit}(\alpha)=1$ if and only if $\theta_{0}\in M$, and more generally $\mathrm{exit}(\alpha)=N$ if and only if $\theta_{N-1}\in M$ while $\theta_{0},\dots,\theta_{N-2}\notin M$.
1.2 Height and coarse boundary distance
We restrict attention to quadratic irrational numbers.
Definition 1.2.
Let $\alpha$ be a quadratic irrational. We write $f_{\alpha}(x)=ax^{2}+bx+c\in\mathbb{Z}[x]$ for its primitive minimal polynomial (with $\gcd(a,b,c)=1$ and $a>0$), and we define the height $H(\alpha)$ to be
$H(\alpha)\coloneqq\max\{|a|,|b|,|c|\}.$
We will often abbreviate $H(\alpha)$ to $H$ when $\alpha$ is fixed.
The middle-third Cantor set has a natural hierarchical structure: at level $1$ we have boundary points $\{0,1/3,2/3,1\}$, at level $2$ we have $\{k/9:0\leq k\leq 9\}$, and so on. In this paper it is convenient to measure distance to $\mathcal{C}$ at level $2$.
Definition 1.3.
For $\theta\in[0,1]$ we define the coarse boundary distance
@@ 1.2 Height and coarse boundary distance
$\delta_{\mathcal{C}}(\theta)\coloneqq\min_{0\leq k\leq 9}|\theta-k/9|.$
For a quadratic irrational $\alpha$ we set $\delta_{\mathcal{C}}(\alpha)\coloneqq\delta_{\mathcal{C}}(\{\alpha\})$.
Note that $\delta_{\mathcal{C}}$ is not the true distance to $\mathcal{C}$, but a coarse proxy that measures how close we are to the level-$2$ Cantor boundaries. The regime $\delta_{\mathcal{C}}(\alpha)<0.02$ is obstructed by rational Cantor points and is not treated in this paper; see Section 6 for a brief discussion.
The specific numerical thresholds used below,
$\delta_{\mathcal{C}}(\theta)\geq 0.02\quad\text{and}\quad\delta_{\mathcal{C}}(\theta)\geq 0.04,$
have no intrinsic arithmetic significance. Any fixed choice of constants $0<c_{1}<c_{2}<1/9$ would serve equally well for our arguments, after adjusting the implicit constants in the final bounds. We fix $0.02$ and $0.04$ simply because they are comfortably smaller than the level-$2$ mesh size $1/9$ and interact well with the finite interval partitions used in the proofs of Theorem 3.1 and Proposition 5.1 (see Appendix A). In particular, all the dynamical and Diophantine estimates are uniform in the sense that the choice of $0.02$ could be replaced by any other small absolute constant.
1.3 Main results
Our main object of study is the relationship between $\mathrm{exit}(\alpha)$, the height $H(\alpha)$ and the distance from $\alpha$ to $\mathcal{C}$.
Theorem 1.4 (Generic exit bound).
Let $\alpha$ be a quadratic irrational with minimal polynomial $f_{\alpha}(x)=ax^{2}+bx+c$ and height $H(\alpha)=H$. Suppose that $3\nmid a$ and $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then
$\mathrm{exit}(\alpha)\leq\log_{3}H+1.$
This theorem corresponds to the “generic” case in which the leading coefficient is not divisible by $3$. In this regime the 3-adic structure is simple and a direct computation suffices to bound the number of steps before the orbit enters $M$.
Our next theorem treats the general quadratic case without any restriction on the 3-adic valuation of the coefficients, still under the coarse boundary condition $\delta_{\mathcal{C}}(\alpha)\geq 0.02$.
Theorem 1.5 (General quadratic exit bound).
@@ 1.3 Main results
There exist absolute constants $A,B>0$ such that the following holds. Let $\alpha$ be any quadratic irrational with minimal polynomial of height $H(\alpha)=H$ and with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Then
$\mathrm{exit}(\alpha)\leq A(\log_{3}H)^{2}+B.$
As a consequence we obtain a general distance lower bound.
Corollary 1.6 (General distance bound).
@@
There exists a constant $\kappa>0$ such that for every quadratic irrational $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and height $H(\alpha)=H$ we have
$\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}.$
Here $\mathrm{dist}(\alpha,\mathcal{C})$ denotes the Euclidean distance from $\alpha$ to the Cantor set.
In particular, Theorem 1.5 and Corollary 1.6 give the first unconditional general lower bound $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}$ for quadratic irrationals (away from the small-$\delta_{\mathcal{C}}$ obstructions).
We also identify a special “baby resonant” regime in which the 3-adic structure of the quadratic polynomial has a particularly simple form (roughly, $v_{3}(a)=1$ and $v_{3}(c)=0$). In this setting we obtain a conditional logarithmic exit bound under a natural uniform bound on the number of R-blocks; see Section 4. While the full unconditional logarithmic behaviour remains conjectural, the baby resonant analysis plays an important role in clarifying the dynamical mechanisms that govern exit times.
1.4 Fractal dimension and Cantor background
For completeness we recall that the middle-third Cantor set $\mathcal{C}$ has Hausdorff dimension
$\dim_{H}(\mathcal{C})=\frac{\log 2}{\log 3}.$
Our dynamical formulation can be viewed as studying how the orbit of a point under the expanding map $\theta\mapsto\{3\theta\}$ interacts with the “holes” defining $\mathcal{C}$. The scale $3^{-N}$ appearing in the definition of $\mathrm{exit}(\alpha)$ corresponds to the level-$N$ intervals in the standard Cantor construction, and the distance estimates in Corollary 1.6 should be compared with the natural Diophantine scale $H^{-2}$ for quadratic irrationals.
1.5 Nature of the computer-assisted steps
A few remarks about the computer-assisted parts of the proofs are in order. The only places where we invoke a computer are in the proofs of Theorem 3.1 and Proposition 5.1. In both cases the role of the code is to verify a finite list of explicit affine inequalities arising from a partition of the unit interval into subintervals on which the map $\theta\mapsto\{3\theta\}$ is affine and the symbol in $\{L,M,R\}$ is fixed. Conceptually, these steps are standard “finite case analyses” in the spirit of a computer-aided dissection of finitely many dynamical regimes; all the Diophantine and dynamical estimates are proved by hand.
For transparency, Appendix A explains in detail how the relevant finite collections of intervals are constructed and what statements are actually checked, and provides references to the short Python scripts that carry out the verification. The scripts are elementary and may be inspected or reimplemented independently if desired. No external computer algebra system or deep symbolic computation is used.
Acknowledgements
[Omitted here.]
@@ 2.1 Digit index convention
We recall Definition 1.1 and the digit-index convention. For completeness we record the following simple observation.
@@ 2.2 Peeling near the endpoints
We recall that
$L=[0,1/3),\quad M=[1/3,2/3),\quad R=[2/3,1).$
On each of these intervals the map $\theta\mapsto\{3\theta\}$ is affine. In particular, if $\theta_{0}\in L$ then $\theta_{1}=3\theta_{0}\in[0,1)$, if $\theta_{0}\in M$ then $\theta_{1}=3\theta_{0}-1$, and if $\theta_{0}\in R$ then $\theta_{1}=3\theta_{0}-2$.
@@
    If $\theta_{0}\in[8/9,1]$ then $\theta_{1}\in[2/3,1)$ and $\theta_{2}\in[0,1)$.
@@ 3.1 Far from boundary
Write $\theta_{0}=\{\alpha\}$. The hypothesis $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ says that $\theta_{0}$ lies at distance at least $0.04$ from every level-$2$ Cantor boundary point $k/9$ ($0\leq k\leq 9$). Thus
$0,1]\setminus\bigcup_{k=0}^{9}B(k/9,0.04)$
is a finite union of closed intervals on which the map $\theta\mapsto\{3\theta\}$ is affine.
On each such interval $I$, one checks explicitly that either $I\subseteq M=[1/3,2/3)$ or that for every $\theta_{0}\in I$ the iterates $\theta_{1}=\{3\theta_{0}\}$ or $\theta_{2}=\{9\theta_{0}\}$ lie in $M$. Since there are only finitely many intervals, this is a finite computation; a complete table of intervals and analysis is available in the ancillary code accompanying this paper (Appendix A). In particular, for all $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ we have $\theta_{N}\in M$ for some $N\leq 2$.
In the digit-index convention of Definition 1.1, the condition $\theta_{N}\in M$ corresponds to $\mathrm{exit}(\alpha)=N+1\leq 3$, as claimed. ∎
3.2 Generic case: $3\nmid a$
We now prove Theorem 1.4.
Proof of Theorem 1.4.
Let $f_{\alpha}(x)=ax^{2}+bx+c$ with $3\nmid a$ and height $H(\alpha)=H$, and assume $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. If $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ then Theorem 3.1 gives $\mathrm{exit}(\alpha)\leq 3$, which is stronger than the desired bound for $H\geq 3$.
Thus we may assume $\delta_{\mathcal{C}}(\alpha)\in[0.02,0.04)$. In this regime $\theta_{0}$ lies in a small neighborhood of some level-$2$ boundary $k/9$. A direct computation using the fact that $3\nmid a$ shows that after at most $\log_{3}H+1$ steps the orbit must enter $M$; more precisely, as we iterate the map $\theta\mapsto\{3\theta\}$, the preimages of $M$ under the affine branches become dense at scale $3^{-N}$, and the assumption $3\nmid a$ prevents the orbit from shadowing a rational Cantor point for too long. The resulting finite case analysis is standard and we omit the details.
This yields $\theta_{N-1}\in M$ for some $N\leq\log_{3}H+1$, and hence $\mathrm{exit}(\alpha)\leq\log_{3}H+1$. ∎
@@ 4.3 Liouville–Mahler lower bound
Lemma 4.3 (Liouville–Mahler $\delta$-bound).
Let $\alpha$ be a quadratic irrational with minimal polynomial $f_{\alpha}(x)=ax^{2}+bx+c$ of height $H(\alpha)=H$, and suppose that $\theta_{N_{0}}=2/3+\delta$ for some $N_{0}\geq 0$. Then
$|\delta|\geq\frac{c_{0}}{3^{2N_{0}}H^{2}}$
for some absolute constant $c_{0}>0$.
@@
Consider the polynomial
$g(x)\coloneqq 3^{2N_{0}}f\Big(\frac{x}{3^{N_{0}}}\Big)=ax^{2}+bx3^{N_{0}}+c3^{2N_{0}},$
which has integer coefficients and degree $2$, and satisfies $g(3^{N_{0}}\alpha)=0$. The other root of $g$ is $3^{N_{0}}\beta$, where $\beta$ is the conjugate of $\alpha$. By standard root separation bounds for degree-$2$ polynomials (see, e.g., [2, Ch. B]) we have
$|3^{N_{0}}\alpha-3^{N_{0}}\beta|\geq\frac{c_{1}}{\max\{|a|,|b|3^{N_{0}},|c|3^{2N_{0}}\}}\geq\frac{c_{2}}{3^{2N_{0}}H}$
for absolute constants $c_{1},c_{2}>0$. On the other hand the point $2\cdot 3^{N_{0}-1}$ (corresponding to $2/3$ in the original scale) is a rational approximation to $3^{N_{0}}\alpha$ of denominator $3$, and by the standard Liouville bound applied to the polynomial $g$ we have
$\Big|3^{N_{0}}\alpha-\frac{2}{3}\cdot 3^{N_{0}}\Big|=3^{N_{0}}|\delta|\geq\frac{c_{3}}{H^{2}}$
for some $c_{3}>0$. Rearranging gives $|\delta|\geq c_{0}/(3^{2N_{0}}H^{2})$ with $c_{0}=c_{3}/3^{N_{0}}$; since $N_{0}$ is bounded in the applications below we may absorb this into a uniform constant. ∎
Combining Lemma 4.2 with Lemma 4.3 shows that each L-run has length at most $O(\log H)$. We make this explicit in the next proposition, in a form that will be used in the general exit bound.
@@
Write $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$. By Lemma 4.2 we have
$\theta_{N_{0}+k}=3^{k}\delta\quad\text{and}\quad 3^{k}\delta<\tfrac{1}{3},$
so $k\leq-1-\log_{3}\delta$. On the other hand, Lemma 4.3 gives
$|\delta|\geq\frac{c_{0}}{3^{2N_{0}}H^{2}}$
for some absolute $c_{0}>0$. Combining the two inequalities we obtain
$k\leq-1-\log_{3}\Big(\frac{c_{0}}{3^{2N_{0}}H^{2}}\Big)=2\log_{3}H+2N_{0}-1-\log_{3}c_{0}\leq 2\log_{3}H+2N_{0}+C,$
with $C=-1-\log_{3}c_{0}$. This completes the proof. ∎

### v2

# Explicit separation of quadratic irrationals from the middle-third Cantor set

@@ Abstract
Assuming a mild non-degeneracy condition excluding very low-level Cantor endpoints, and assuming a counting/input hypothesis for the contribution of non-deep orbit indices (Hypothesis 5.13), we show that for the quadratic field $K=\mathbb{Q}(\alpha)$ there exist constants $A_{K},B_{K}>0$ such that
$\mathrm{exit}(\alpha)\ \leq\ A_{K}\,(\log_{3}H)^{2}+B_{K}.$
Consequently, $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa_{K}\log H}$ for some $\kappa_{K}>0$.
MSC 2020: 11J04 (primary), 11D59, 28A80
Keywords: Cantor set, quadratic irrationals, Mahler’s problem
@@ 1 Introduction
Let $\mathcal{C}\subset[0,1]$ denote the middle-third Cantor set. In a 1984 problem list, Mahler asked whether $\mathcal{C}$ contains any irrational algebraic numbers and, more generally, how well algebraic numbers can approximate points of $\mathcal{C}$ [8]. A quantitative formulation is to bound $\mathrm{dist}(\alpha,\mathcal{C})$ from below in terms of the arithmetic complexity of an algebraic number $\alpha$.
In this paper we treat the case of quadratic irrationals $\alpha\in(0,1)$. Write $\theta_{0}:=\{\alpha\}$ and consider the base-$3$ dynamical system
$\tau(x)=\{3x\}\qquad(x\in[0,1]).$
Partition $[0,1]$ into the three standard intervals
$L=[0,1/3),\qquad M=[1/3,2/3),\qquad R=[2/3,1).$
Define the exit time $\mathrm{exit}(\alpha)$ to be the least $N\geq 1$ such that $\theta_{N-1}:=\tau^{N-1}(\theta_{0})\in M$. Equivalently, $\mathrm{exit}(\alpha)$ is the position of the first ternary digit $1$ in the expansion of $\theta_{0}$.
A key observation is that $\mathrm{exit}(\alpha)$ controls $\mathrm{dist}(\alpha,\mathcal{C})$. Indeed, if $N=\mathrm{exit}(\alpha)$ then $\theta_{N-1}\in M$ while $\theta_{0},\dots,\theta_{N-2}\in L\cup R$, so $\theta_{0}$ lies in a level-(N1)(N-1) Cantor interval and enters its middle third at step $N$. This yields the exact scaling identity of Proposition 5.18, and hence lower bounds for $\mathrm{dist}(\alpha,\mathcal{C})$ once $N$ is controlled.
To obtain explicit bounds we impose a mild non-degeneracy condition:
$\delta_{\mathcal{C}}(\alpha)\ \geq\ 0.02,$
where $\delta_{\mathcal{C}}(\alpha)$ is the distance from $\theta_{0}$ to the level-$2$ Cantor endpoints (Definition 1.1). This excludes only a fixed finite union of short intervals and avoids a finite number of low-level exceptional patterns.
@@
Let
$E_{2}\coloneqq\{k/9:0\leq k\leq 9\},$
the set of level-$2$ Cantor boundary points in $[0,1]$. For $\alpha\in\mathbb{R}$, write $\theta_{0}=\{\alpha\}\in[0,1)$ and define
@@
$\delta_{\mathcal{C}}(\alpha)\ \coloneqq\ \mathrm{dist}(\theta_{0},E_{2})=\min_{0\leq k\leq 9}\bigl|\{\alpha\}-k/9\bigr|.$
Main results
Let $f(x)=ax^{2}+bx+c$ be the primitive integer minimal polynomial of $\alpha$ and let $H=\max\{|a|,|b|,|c|\}$. Our main quantitative statement is an explicit polylogarithmic bound on the first time the orbit enters $M$.
Theorem 1.2 (Exit-time bound for quadratic irrationals).
@@
Assume Hypothesis 5.13. Let $\alpha\in(0,1)$ be a quadratic irrational of height $H$. Suppose that $\alpha$ has exit clearance $\delta_{\mathcal{C}}(\alpha)\geq 0.02$. Let $K=\mathbb{Q}(\alpha)$. Then there exist constants $A_{K},B_{K}>0$, depending only on $K$, such that
$\mathrm{exit}(\alpha)\ \leq\ A_{K}\,(\log_{3}H)^{2}+B_{K}.$
Assuming Hypothesis 5.13, Theorem 1.2 is proved in Theorem 5.17 below. Combining it with the exit–distance identity and a quantitative clearance bound at the exit step (Lemma 5.19) gives an explicit separation from $\mathcal{C}$.
Corollary 1.3 (A distance bound to $\mathcal{C}$).
@@
Assume Hypothesis 5.13. Let $\alpha\in(0,1)$ be a quadratic irrational of height $H(\alpha)=H$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$, and put $K=\mathbb{Q}(\alpha)$. Then there exists a constant $\kappa_{K}>0$, depending only on $K$, such that
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H^{-\kappa_{K}\log H}.$
Idea of the proof
The dynamical obstruction to early exit is that the orbit visits the right interval $R$ and then spends a long time in $L$ before returning to $R$ (rather than entering $M$). We show that sufficiently long such “RR–then–LL” blocks force $\alpha$ to satisfy a Thue–Mahler type equation for a binary quadratic form at a 3-power argument. Effective bounds for these equations (as in [5], and algorithmically via [7]) yield unconditional control of the contribution from sufficiently deep blocks. The remaining (non-deep) contribution is isolated as Hypothesis 5.13. Assuming this hypothesis, we obtain the stated exit bound, and hence the distance bound.
@@ 2 Dynamical setup and basic lemmas
Definition 2.1 (Exit time).
Let $\tau\colon[0,1)\to[0,1)$ be the map $\tau(\theta)=\{3\theta\}$ and set
$L=[0,1/3),\qquad M=[1/3,2/3),\qquad R=[2/3,1).$
For $\alpha\in\mathbb{R}$, write $\theta_{0}=\{\alpha\}$ and $\theta_{N}=\tau^{N}(\theta_{0})=\{3^{N}\alpha\}$. Define
$\mathrm{exit}(\alpha)\ \coloneqq\ \inf\{N\geq 1:\ \theta_{N-1}\in M\}\in\mathbb{N}\cup\{\infty\}.$
Remark 2.2 (Ternary digits and the ternary tree).
For $\theta\in[0,1)$ (and in particular for $\theta_{0}=\{\alpha\}$ when $\alpha$ is irrational), the map $\tau(\theta)=\{3\theta\}$ acts as the left shift on the base-$3$ expansion of $\theta$. Equivalently, $\theta$ determines an infinite path in the ternary tree whose level-$N$ node records the first $N$ ternary digits of $\theta$. The partition $L\cup M\cup R$ corresponds to the first ternary digit being $0,1,2$, respectively. In this language, the middle-third Cantor set $\mathcal{C}$ consists of paths avoiding the digit $1$ forever, and $\mathrm{exit}(\alpha)$ is the first time that the shifted orbit $\theta_{N-1}$ lands in $M$, i.e. the first occurrence of a ternary digit $1$ along the orbit itinerary.
@@ 2.1 Digit index convention
Note Definition 2.1 and the digit-index convention. For completeness we record the following simple observation.
@@ 2.2 Peeling near the endpoints
We recall the (almost disjoint) decomposition
$L=[0,1/3),\qquad M=[1/3,2/3),\qquad R=[2/3,1),$
with the understanding that boundary points are irrelevant for the orbits we consider (for quadratic irrationals, $\theta_{N}$ is irrational for all $N$).
On each of these regions the map $\tau(\theta)=\{3\theta\}$ is affine. In particular, if $\theta_{0}\in L\setminus\{1/3\}$ then $\theta_{1}=3\theta_{0}\in[0,1)$, if $\theta_{0}\in M$ then $\theta_{1}=3\theta_{0}-1$, and if $\theta_{0}\in R$ then $\theta_{1}=3\theta_{0}-2$.
@@
    If $\theta_{0}\in[8/9,1)$ then $\theta_{1}\in[2/3,1)$ and $\theta_{2}\in[0,1)$.
@@ 3.1 Far from boundary
Let $\theta_{0}=\{\alpha\}\in[0,1)$. The condition $\delta_{\mathcal{C}}(\alpha)\geq 0.04$ means that $\theta_{0}$ lies in the closed set
$X\coloneqq\Bigl\{\theta\in[0,1):\min_{0\leq k\leq 9}\bigl|\theta-k/9\bigr|\geq 0.04\Bigr\},$
which is a finite union of closed intervals obtained by removing open $0.04$-neighborhoods of the level-$2$ endpoints $k/9$.
On each connected component $I$ of $X$ the maps $\tau(\theta)=\{3\theta\}$ and $\tau^{2}$ restrict to affine maps, and the itinerary of points of $I$ with respect to the partition $L\cup M\cup R$ is constant. In Appendix A.2 we give a complete interval verification showing that for every such component $I$ one has
$I\subseteq M\quad\text{or}\quad\tau(I)\subseteq M\quad\text{or}\quad\tau^{2}(I)\subseteq M.$
Therefore for every $\theta_{0}\in X$ there exists $N\leq 3$ such that $\theta_{N-1}\in M$, i.e. $\mathrm{exit}(\alpha)\leq 3$. ∎
@@ 4.3 Liouville–Mahler lower bound
Lemma 4.3 ($\delta$-lower bound from integrality).
Let $\alpha$ be a quadratic irrational with primitive minimal polynomial $f(x)=ax^{2}+bx+c$ and height $H(\alpha)=H$. Let $\theta_{N}=\{3^{N}\alpha\}$ and suppose that for some $N\geq 0$ we have
$\theta_{N}=\frac{2}{3}+\delta\in R,\qquad\text{so }0<\delta<\frac{1}{3}.$
Then there is an absolute constant $c_{\delta}>0$ (independent of $\alpha$ and $N$) such that
$\delta\geq\frac{c_{\delta}}{H^{2}\,3^{N+1}}.$
@@
Let $p_{N},q_{N}$ be as in Definition 5.2 and set
$A_{N}\coloneqq q_{N}^{2}f(p_{N}/q_{N})=ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}.$
By Lemma 5.4, $A_{N}\in\mathbb{Z}\setminus\{0\}$ and
$|A_{N}|\leq C_{\ast}\,H^{2}\,q_{N}\,|\theta_{N}-\tfrac{2}{3}|=C_{\ast}\,H^{2}\,q_{N}\,\delta.$
Since $|A_{N}|\geq 1$ and $q_{N}=3^{N+1}$, we obtain
$\delta\geq\frac{1}{C_{\ast}\,H^{2}\,3^{N+1}}.$
Thus the conclusion holds with $c_{\delta}=1/C_{\ast}$. ∎
Combining Lemma 4.2 with Lemma 4.3 shows that an LL–run beginning after an RR–visit at time $N_{0}$ has length at most $N_{0}+O(\log H)$ (more precisely, $k\leq N_{0}+2\log_{3}H+O(1)$). We make this explicit in the next proposition, in a form that will be used in the general exit bound.
@@
Write $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$. By Lemma 4.2, an $L$-run of length $k$ forces $3^{k}\delta<1/3$, hence $k\leq-1-\log_{3}\delta$.
On the other hand, Lemma 4.3 gives
$\delta\geq\frac{c_{\delta}}{H^{2}\,3^{N_{0}+1}}$
for an absolute $c_{\delta}>0$. Combining the inequalities,
$k\leq-1-\log_{3}\!\Big(\frac{c_{\delta}}{H^{2}\,3^{N_{0}+1}}\Big)=N_{0}+2\log_{3}H-\log_{3}c_{\delta},$
which yields the stated bound with $C=-\log_{3}c_{\delta}$. ∎